A 8.00L tank at 17.6°C is filled with 3.25g of sulfur tetrafluoride gas and 12.2g of chlorine pentafluoride gas. You can assume both gases behave as ideal gases under these conditions.
Calculate the mole fraction and partial pressure of each gas, and the total pressure in the tank. Be sure your answers have the correct number of significant digits.
1
Expert's answer
2020-04-27T01:45:05-0400

pV=nRT. p=nRT/V= mRT/M= 3.25 x 8.314 x 290.6 / 108 = 72.7 kpa.

p= 12.2 x 8.314 x 290.6 / 130.5 = 225.86 kpa.

ptotal=298.56 kpa.

Mole fraction of sulfur tetrafluoride 72.7/298.56 = 0.24.

and 1-0.24 = 0.76.